# Scalability Challenges in Variational Quantum Optimization under Stochastic Noise

### v1

@@ Appendix A Mapping QUBO to Ising Problems
A similar mapping between the QUBO and the Maximum Cut (MaxCut) problem was derived in Ref. [110].
@@ Appendix B Quadratic Penalty Terms in Permutation Problems
In graph partitioning problems, such as the well-known MaxCut problem, the number of nodes in the problem graph corresponds directly to the number of binary variables in the problem formulation. However, permutation problems, such as the highly relevant Traveling Salesperson Problem (TSP), are more complex, as they require assigning an order instead of a partition to the nodes. This can be achieved using one-hot-encoding, where permutations are represented by $n^{2}$ binary decision variables $x_{i}^{\alpha}$, indicating whether node $i$ is assigned to a position $\alpha$ in the permutation. In the case of the TSP, $i$ would represent a specific city, and $\alpha$ denotes the time step at which the city is visited along the route.
@@
Here, $C_{0}(\mathbf{x})$ represents the actual unconstrained loss function of the problem, defined by the QUBO matrix $\mathbf{Q_{0}}$, while the latter terms impose the corresponding permutational constraints. In the TSP, $C_{0}(\mathbf{x})$ would be a function that returns the total path length.
To ensure that feasible solutions are enforced after minimizing Eq. 45, $P$ must be larger than a problem-specific threshold depending on $C_{0}$. To obtain the QUBO matrix representation from the loss function in Eq. 45, we translate its terms into the matrix elements $Q_{ij}^{\alpha\beta}$, which encode the interaction between variables $x_{i\alpha}$ and $x_{j\beta}$:
@@
From Eq. 46, it is immediately apparent that the penalty terms affect the diagonal and off-diagonal elements of $\mathbf{Q}$ differently:
@@
Since $P$ is chosen to be significantly larger than each element of $\mathbf{Q_{0}}$, all diagonal elements of $\mathbf{Q}$ become negative. This confirms that adding equality constraints as quadratic penalty terms transforms the problem’s QUBO matrix into one with negative diagonal entries, justifying the formulation used in Section III.1.
@@ Appendix C Distribution of Finite Sampling Errors
As quantum computing advances toward fully error-corrected systems, the only remaining source of stochastic noise will be that arising from the finite sampling of circuit measurements. In the VQA paradigm (cf. Section II.2), the number of measurement shots $n_{\mathrm{shots}}$ directly determines the precision of the estimated expectation value $\mathbb{E}[\mathbf{C}]$ relative to the true loss value $\langle\mathbf{C}\rangle$.
@@
Here, $p_{j}=\lvert\langle\bm{\psi}(\bm{\theta})|\bm{q}_{j}\rangle\rvert^{2}$ is the probability of measuring the system in the basis state $|\bm{q}_{j}\rangle$, which can only be estimated from measurement samples on a quantum device. The measurement counts of the basis states follow a multinomial distribution such that each estimated proportion $\hat{p}_{j}$ carries the standard error given in Eq. 47.
@@
In Section III.3, we argue that these finite sampling errors can be reasonably modeled by adding Gaussian noise to the exact values of the loss function, assuming that this is a good-enough characterization of how statistical errors are distributed across the parameter space during loss evaluation. To illustrate this, Fig. 14 presents the distribution of total normalized errors $\delta\mathcal{L}$ when sampling the loss at a randomly chosen point $\bm{\theta}\in[-2\pi,2\pi]^{n}$ for a system of size $n=6$ with $n_{\mathrm{shots}}=1024$ shots. These distributions are computed across the three VQAs discussed in Section III.2 and compared to a normal distribution with a mean of zero and the same standard deviation as the empirical data. The close visual agreement between the histogram and the theoretical Gaussian curve validates the central limit theorem, which predicts that for a sufficiently large number of shots, the binomial (or multinomial) distribution from finite sampling converges to a normal distribution.
@@
Figure 14: Distribution of finite sampling errors across loss landscape. (a) Standard deviation of the total errors from $1000$ loss evaluations with $1024$ measurement shots, at $100$ randomly selected points across the loss landscape of an $n=6$ QUBO problem instance (cf. Section III.1). The relationship between the normalized loss value and the error magnitude is analyzed for three different VQAs, with linear fits indicating whether the noise behaves additively or multiplicatively. (b) Probability density histograms of $1000$ normalized errors at a single representative point, compared to the theoretical normal distribution with zero mean and the same standard deviation, validating the central limit theorem.
@@
Fig. 14 also investigates the relationship between the total error magnitude $s[\delta\mathcal{L}]$ and the loss value at each evaluation point. This analysis helps to determine whether finite sampling noise behaved in an additive or multiplicative manner. To quantify this, linear fits of the sample standard deviations are performed, yielding values for the slope $m$ and y-intercept $t$ of the lines: If $m\ll t$, the noise is predominantly additive, whereas if $m\gg t$, the noise exhibits a multiplicative behavior.
@@
For the BENQO algorithm, the y-intercept $t$ is significantly larger than the slope $m$, meaning that the error is almost entirely additive. For the other two algorithms, the error displays both additive and multiplicative components, with a larger multiplicative contribution. Their increased multiplicative component arises from the strong dependence of the statistical error (49) on $\langle\mathbf{C}\rangle$. BENQO, however, almost entirely loses this dependency due to its implicit measurement strategy [45].
@@
Despite these differences, additive errors tend to distort the loss landscape more severely than multiplicative ones [52]. Therefore, when modeling total finite sampling-induced errors, a worst-case scenario can always be represented by simply adding Gaussian noise to the loss function. This justifies the noise model used in our methodology (cf. Section III.3).
@@ Appendix D Ranking of Tested Optimizers
Certain optimizers – CG, SLSQP, and BFGS – were excluded from this ranking due to their poor performance across all tested noise levels. Their failure likely stems from their reliance on precise gradient calculations, which become very unreliable in the presence of noise. This supports previous findings [75, 74], indicating that gradient-based optimizers, while efficient under ideal conditions, struggle in noisy settings, whereas gradient-free strategies prove significantly more robust.
Lastly, note that while this ranking provides a comparative assessment, it is important to recognize its limitations. The approach, inspired by Singh et al. [74], assigns discrete ranks rather than directly plotting continuous performance metrics. As a result, it does not fully capture quantitative performance differences between optimizers. A more precise analysis would involve directly plotting the number of function calls $n_{\text{calls}}$ against the noise resilience threshold $\sigma^{*}$ (27), allowing for a more detailed examination of performance trade-offs.
@@ Appendix E Optimality Distribution over Classical Solution Space
The upper panel of Fig. 16 presents the total percentage of solutions $p_{t}$ that achieve a loss value within different relative thresholds $t$ of the optimal loss (cf. Eq. 23) across all $2^{n}$ possible classical solution states, i.e., bit strings. Note that this corresponds to the success probabilities of random guessing. Interestingly, this percentage appears to converge for system sizes beyond $7$ or $8$. More importantly, nearly $20\%$ of all possible solution states still yield loss values within $90\%$ of the optimum, indicating that the solution space is highly structured rather than uniformly random.
The lower panel of Fig. 16 shows the corresponding normalized loss distribution of all solution states, which also seems to converge to a fixed shape, Especially, as system size grows, this distribution exhibits a left-skewed shape, with a higher concentration of solutions near the minimum energy. This deviation from a normal distribution explains why solutions above high $t$ values are more common than one might expect from a purely random search. This tendency of solutions to cluster closer to the global minimum than to the global maximum directly influences the behavior of $\hat{p}_{99\%}$, and $\hat{p}_{95\%}$ in our experiments.
Figure 16: Distribution of solution optimality. (a) The percentage of solutions $p_{t}$ that achieve a loss value within a relative threshold $t$ of the optimal loss value across varying system sizes $n$ is depicted as contours. (b) Violin plots illustrate the distribution of all $2^{n}$ possible classical solution states within the normalized loss landscape $\hat{\mathcal{L}}$. The loss value corresponding to the $95\%$ optimality threshold is marked for each $n$, providing insight into the spread and concentration of near-optimal solutions.
@@ Appendix F Fit Analysis and Parameters for Observed Noise Resilience Decay
Fig. 8 presents the fitted decay functions (cf. Eqs. 30, 31 and 32) for the declining noise resilience threshold $\sigma^{*}(n)$ across all tested optimizers. To assess the quality of these fits, we compute the mean squared error (MSE)
@@
NGD NFT COBYLA Powell SPSA
@@
Table 2: MSE of decay function fits. The MSE values for each optimizer’s noise resilience threshold $\sigma^{*}(n)$ are computed for three decay models, exponential (exp), power-law (pl), and logarithmic (log), following Eqs. 30, 31 and 32. The lowest MSE for each optimizer is marked for better comparability.
NGD NFT COBYLA Powell SPSA
@@
Table 3: Final parameters of decay function fits. The optimized parameters $k^{*}$ and $\gamma^{*}$ for the exponential (exp), the power-law (pl) and logarithmic (log) decay functions (cf. Eqs. 30, 31 and 32) are listed for the measured noise resilience thresholds $\sigma^{*}(n)$ (27) and $\epsilon^{*}(n)$ (35). Results are provided for all tested optimizers based on Fig. 8.

### v2

@@ Appendix A Mapping QUBO to Ising Problems
A similar mapping between the QUBO and Maximum Cut (MaxCut) problem was derived in Ref. [134].
@@ Appendix B Quadratic Penalty Terms in Permutation Problems
In graph partitioning problems, such as the well-known MaxCut problem, the number of nodes in the problem graph directly corresponds to the number of binary variables in the problem formulation. However, permutation problems, such as the highly relevant Traveling Salesperson Problem (TSP), are more complex as they require assigning an order instead of a partition to the nodes. This can be achieved using one-hot-encoding, where permutations are represented by $n^{2}$ binary decision variables $x_{i}^{\alpha}$, indicating whether node $i$ is assigned to a position $\alpha$ in the permutation. In the case of the TSP, $i$ would represent a specific city, and $\alpha$ denotes the time step at which the city is visited along the route.
@@
Here, $C_{0}(\mathbf{x})$ represents the actual unconstrained loss function of the problem, defined by the QUBO matrix $\mathbf{Q_{0}}$, while the latter terms impose the corresponding permutational constraints. In the TSP, $C_{0}(\mathbf{x})$ would be a function returning the total path length.
To ensure that feasible solutions are enforced after minimizing Eq. 45, $P$ must be larger than a problem-specific threshold dependent on $C_{0}$. To obtain the QUBO matrix representation from the loss function in Eq. 45, we translate its terms into the matrix elements $Q_{ij}^{\alpha\beta}$, which encode the interaction between variables $x_{i\alpha}$ and $x_{j\beta}$:
@@
From Eq. 46, it is immediately apparent that the penalties terms affect the diagonal and off-diagonal elements of $\mathbf{Q}$ differently:
@@
Since $P$ is chosen to be significantly larger than each element of $\mathbf{Q_{0}}$, all diagonal elements of $\mathbf{Q}$ become negative. This confirms that adding equality constraints as quadratic penalty terms transforms the problem’s QUBO matrix into one with negative diagonal entries, justifying the formulation used in Section IV.1.
@@ Appendix C Distribution of Finite Sampling Errors
As quantum computing advances toward fully error-corrected systems, the only remaining source of stochastic noise will be that arising from the finite sampling of circuit measurements. In the VQA paradigm (cf. Section III.2), the number of measurement shots $n_{\mathrm{shots}}$ directly determines the precision of the estimated expectation value $\mathbb{E}[\mathbf{C}]$ relative to the true loss value $\langle\mathbf{C}\rangle$.
@@
Here, $p_{j}=\lvert\langle\bm{\psi}(\bm{\theta})|\bm{q}_{j}\rangle\rvert^{2}$ is the probability of measuring the system in the basis state $|\bm{q}_{j}\rangle$, which can only be estimated from measurement samples on a quantum device. The measurement counts of basis states follow a multinomial distribution, such that each estimated proportion $\hat{p}_{j}$ carries the standard error given in Eq. 47.
@@
In Section IV.3, we argue that these finite sampling errors can be reasonably modeled by adding Gaussian noise to the exact values of the loss function, assuming that this is a good-enough characterization of how statistical errors are distributed across the parameter space during loss evaluation. To illustrate this, Fig. 14 presents the distribution of total normalized errors $\delta\mathcal{L}$ when sampling the loss at a randomly chosen point $\bm{\theta}\in[-2\pi,2\pi]^{n}$ for a system of size $n=6$ with $n_{\mathrm{shots}}=1024$ shots. These distributions are computed across the three VQAs discussed in Section IV.2 and compared to a normal distribution with a mean of zero and the same standard deviation as the empirical data. The close visual agreement between the histogram and the theoretical Gaussian curve validates the central limit theorem, which predicts that for a sufficiently large number of shots, the binomial (or multinomial) distribution from finite sampling converges to a normal distribution.
@@
Figure 14: Distribution of finite sampling errors across loss landscape. (a) Standard deviation of the total errors from $1000$ loss evaluations with $1024$ measurement shots, at $100$ randomly selected points across the loss landscape of an $n=6$ QUBO problem instance (cf. Section IV.1). The relationship between the normalized loss value and the error magnitude is analyzed for three different VQAs, with linear fits indicating whether the noise behaves additively or multiplicatively. (b) Probability density histograms of $1000$ normalized errors at a single representative point, compared to the theoreticcal normal distribution with zero mean and the same standard deviation, validating the central limit theorem.
@@
Fig. 14 also investigates the relationship between the total error magnitude $s[\delta\mathcal{L}]$ and the loss value at each evaluation point. This analysis helps determine whether finite sampling noise behaved in an additive or multiplicative manner. To quantify this, linear fits of the sample standard deviations performed, yielding values for the slope $m$ and y-intercept $t$ of the lines: If $m\ll t$, the noise is predominantly additive, whereas if $m\gg t$, the noise exhibits a multiplicative behavior.
@@
For the BENQO algorithm, the y-intercept $t$ is significantly larger than the slope $m$, meaning that the error is almost entirely additive. For the other two algorithms, the error displays both additive and multiplicative components, with a larger multiplicative contribution. Their increased multiplicative component arises from the strong dependence of the statistical error (49) on $\langle\mathbf{C}\rangle$. BENQO, however, almost entirely looses this dependency due to its implicit measurement strategy [64].
@@
Despite these differences, additive errors tend to distort the loss landscape more severely than multiplicative ones [39]. Therefore, when modeling total finite sampling-induced errors, a worst-case scenario can always be represented by simply adding Gaussian noise to the loss function. This justifies the noise model used in our methodology (cf. Section IV.3).
@@ Appendix D Ranking of Tested Optimizers
Certain optimizers – CG, SLSQP, and BFGS – were excluded from this ranking due to their poor performance across all tested noise levels. Their failure likely stems from their reliance on precise gradient calculations, which become very unreliable in the presence of noise. This supports previous findings [33, 35], indicating that gradient-based optimizers, while efficient in ideal conditions, struggle in noisy settings, whereas gradient-free strategies prove significantly more robust.
Lastly, note that while this ranking provides a comparative assessment, it is important to acknowledge its limitations. The approach, inspired by Singh et al. [35], assigns discrete ranks rather than directly plotting continuous performance metrics. As a result, it does not fully capture quantitative performance differences between optimizers. A more precise analysis would involve directly plotting the number of function calls $n_{\text{calls}}$ against the noise resilience threshold $\sigma^{*}$ (27), allowing for a more detailed examination of the performance trade-offs.
@@ Appendix E Optimality Distribution over Classical Solution Space
The upper panel of Fig. 16 presents the total percentage of solutions $p_{t}$ that achieve a loss value within different relative thresholds $t$ of the optimal loss (cf. Eq. 23) across all $2^{n}$ possible classical solution states, i.e., bit strings. Note that this corresponds to the success probabilities of random guessing. Interestingly, this percentage appears to converge for system sizes beyond $7$ or $8$. More importantly, nearly $20\%$ of all possible solution states still yield loss value within $90\%$ of the optimum, indicating that the solution space is highly structured rather than uniformly random.
The lower panel of Fig. 16 shows the corresponding normalized loss distribution of all solution state, which also seems to converge to a fixed shape, Especially, as system size grows, this distribution exhibits a left-skewed shape, with a higher concentration of solutions near the minimum energy. This deviation from a normal distribution explains why solutions above high $t$ values are more common than one might expect from a purely random search. This tendency of solutions to cluster closer to the global minimum than to the global maximum directly influences the behavior of $\hat{p}_{99\%}$, and $\hat{p}_{95\%}$ in our experiments.
Figure 16: Distribution of solution optimality. (a) The percentage of solutions $p_{t}$ that achieve a loss value within a relative threshold $t$ of the optimal loss value, across varying system sizes $n$, is depicted as contours. (b) Violin plots illustrate the distribution of all $2^{n}$ possible classical solution states within the normalized loss landscape $\hat{\mathcal{L}}$. The loss value corresponding to the $95\%$ optimality threshold is marked for each $n$, providing insight into the spread and concentration of near-optimal solutions.
@@ Appendix F Fit Analysis and Parameters for Observed Noise Resilience Decay
Fig. 8 presents the fitted decay functions (cf. Eqs. 30, 31 and 32) for the declining noise resilience threshold $\sigma^{*}(n)$ with respect to $\hat{p}_{\mathrm{opt}}$ across all tested optimizers. To assess the quality of these fits, we compute the mean squared error (MSE)
@@
$\hat{p}_{\mathrm{opt}}$ NGD NFT COBYLA Powell SPSA
@@
$\hat{p}_{95\%}$ NGD NFT COBYLA Powell SPSA
[…]
log
[…]
Table 2: MSE of decay function fits. The MSE values for each optimizer’s noise resilience threshold $\sigma^{*}(n)$ w.r.t. both $\hat{p}_{\mathrm{opt}}$ (upper) and $\hat{p}_{95\%}$ (lower) are computed for three decay models, exponential (exp), power-law (pl), and logarithmic (log), following Eqs. 30, 31 and 32. The lowest MSE for each optimizer is marked for better comparability.
for $\hat{p}_{\mathrm{opt}}$ NGD NFT COBYLA Powell SPSA
@@
for $\hat{p}_{95\%}$ NGD NFT COBYLA Powell SPSA
$k^{*}$ $\gamma^{*}$ $k^{*}$ $\gamma^{*}$ $k^{*}$ $\gamma^{*}$ $k^{*}$ $\gamma^{*}$ $k^{*}$ $\gamma^{*}$
exp $\sigma^{*}$ $1.4(3)$ $0.30(6)$ $5(2)$ $0.6(1)$ $0.8(2)$ $0.73(7)$ $1.1(2)$ $0.52(6)$ $1.2(4)$ $0.5(1)$
$\epsilon^{*}$ $2.9(6)$ $0.13(4)$ $7(2)$ $0.28(6)$ $1.3(2)$ $0.48(4)$ $2.0(3)$ $0.30(4)$ $1.7(5)$ $0.19(6)$
pl $\sigma^{*}$ $3.1(7)$ $1.5(2)$ $12(3)$ $2.3(2)$ $2.1(2)$ $2.89(9)$ $2.8(4)$ $2.2(1)$ $2.8(9)$ $2.0(2)$
$\epsilon^{*}$ $5(1)$ $0.7(1)$ $15(4)$ $1.4(2)$ $3.2(2)$ $2.12(5)$ $4.4(5)$ $1.48(9)$ $3(1)$ $1.1(2)$
log $\sigma^{*}$ $0.77(5)$ $2.2(2)$ $1.27(7)$ $3.3(2)$ $0.130(2)$ $3.99(7)$ $0.34(1)$ $3.2(1)$ $0.41(3)$ $2.9(2)$
$\epsilon^{*}$ $2.5(2)$ $1.2(2)$ $4.0(3)$ $2.2(2)$ $0.43(1)$ $3.06(9)$ $1.09(4)$ $2.2(1)$ $1.3(1)$ $1.7(3)$
Table 3: Final parameters of decay function fits. The optimized parameters $k^{*}$ and $\gamma^{*}$ for the exponential (exp), the power-law (pl) and logarithmic (log) decay functions (cf. Eqs. 30, 31 and 32) are listed for the measured noise resilience thresholds $\sigma^{*}(n)$ (27) and $\epsilon^{*}(n)$ (35). Results are provided for all tested optimizers and both solvability measures $\hat{p}_{\mathrm{opt}}$ (upper part) based on Fig. 8 and $\hat{p}_{95\%}$ (lower part).